Sam I Am is filling his pool at a rate of 3.2 ft per minute. At the same time, Peter Piper has 16 ft. of water in his pool and is emptying it at a rate of 1.6 ft. per minute. When will they have the same amount of water?
Answer by ptaylor(2052)

Let x=number of minutes that elapses before they each have the same amount of water in their pools
After x minutes, SamIAm has 3.2x feet of water
After x minutes, Peter has 16-1.6x feet of water
So, when 3.2x equals 16-1.6x, they will both have the same amount of water
Our equation to solve then is:
3.2x=16-1.6x add 1.6x to each side
3.2x+1.6x=16
4.8x=16
x=3 1/3 or 10/3 min
CK
in 10/3 min, SamIAm has (10/3)*3.2= 10 2/3 feet of water
in 10/3 min, Peter has 16-(10/3)*1.6=16-5 1/3 =10 2/3 feet of water

Hope this helps---ptaylor